jhonyy9 (jhonyy9):

is this limit solved right,correct ? lg mean logarithm base ten lim (lg 100^(x^2) -(lg 10^(x)) = x->1 = lim (lg 10^(2x^2) -(lg 10^(x)) = x->1 = lim (lg (10^(2x^2))/(10^(x)) = x->1 = lim (lg 10^(2x^2 -x)) = x->1 = lim ((2x^2 -x)lg 10) = x->1 = lim (2x^2 -x) = 1 x->1 @TheSmartOne @ganeshie8 @Kainui @Nnesha ty.

jhonyy9 (jhonyy9):

@phi please ,do you agree with this ? ty.

OpenStudy (holsteremission):

Correct, but you could eliminate some steps. \[\begin{align*}\lim_{x\to1}\left(\log_{10}10^{2x^2}-\log_{10}10^x\right)&=\lim_{x\to1}\left(2x^2\log_{10}10-x\log_{10}10\right)\\[1ex] &=\lim_{x\to1}\left(2x^2-x\right)\end{align*}\]No need to make use of all those other properties of logarithms, but what you've done is certainly valid.

OpenStudy (ageta):

jhonny why are you level79

jhonyy9 (jhonyy9):

i m level 74 than i see it now right

jhonyy9 (jhonyy9):

@ageta where you see level 79 ?

OpenStudy (ageta):

http://prntscr.com/cii66v

OpenStudy (ageta):

are you the real jhonyy

jhonyy9 (jhonyy9):

yes sure from Central Europe ,country Hungary,city Budapest with birth country Romania,the part of Transylvania so why ? there are again so much jhonyy9 ?

jhonyy9 (jhonyy9):

@ageta but than you check my profile i m level 74 - i dont know where you see this level 79 - ? sorry

OpenStudy (ageta):

and jhonyy has lots of fan and you dont

jhonyy9 (jhonyy9):

i m jhonny9

jhonyy9 (jhonyy9):

@jhonyy9 check here my profile

jhonyy9 (jhonyy9):

yes sure its possible that you remember on me when i was champion with level 99"s but my this profile was deleted like these so much other again - hope you know about this cases from the month of may this year

jhonyy9 (jhonyy9):

so and i was necessary rebuild a new profile started from level 15

jhonyy9 (jhonyy9):

@jhonyy who are this jhonyy ?

jhonyy9 (jhonyy9):

what you posted there on this ageta Human Calculator Best Response Medals 0 http://prntscr.com/cii66v so this is my posted photo and profile but level 74

jhonyy9 (jhonyy9):

do you know anything about this case - or can i help you ,do you like knowing again about me or about my past profile or ???

jhonyy9 (jhonyy9):

@TheSmartOne what you say about this above wrote please ? ty . this mean that there are other jhonny with my profile photo ? i dont believe it

jhonyy9 (jhonyy9):

ty.

Nnesha (nnesha):

:D haha :=))

Nnesha (nnesha):

you forgot to add street # goin to Europe

jhonyy9 (jhonyy9):

really @Nnesha do you like viste me here - ok. i wait you

OpenStudy (learner):

You can just substitute directly (right away), there is nothing wrong with that. \(\color{blue }{\displaystyle \lim_{x\to1} (\log [100^{x^2}] -\log [10^x])= }\) \(\color{blue }{\displaystyle \lim_{x\to1} (\log [100^{\color{red}{1}^2}] -\log [10^\color{red}{1}]) = }\) \(\color{blue }{\displaystyle \lim_{x\to1} (2 -1)= 1. }\)